OpenStudy (anonymous):

If x=(2+^5)^1/2 + (2-^5)^1/2 and y=(2+^5)^1/2 + (2-^5)^1/2 then evaluate x(square) + y (square)

OpenStudy (anonymous):

5 is power of what term ??

OpenStudy (anonymous):

can you please rewrite by using equation option?

OpenStudy (anonymous):

ill write the question in equation form ...pls.. wait

OpenStudy (anonymous):

\[if x=(2+\sqrt{5)}^{1/2} + (2-\sqrt{5})^{1/2}\]

OpenStudy (anonymous):

\[and y=(2+\sqrt{5})^{1/2} - (2-\sqrt{5})^{1/2}\]

OpenStudy (anonymous):

evaluate \[x ^{2}+y ^{2}\]

OpenStudy (anonymous):

nw can u help me find the answer

OpenStudy (shubhamsrg):

note that xy = 2sqrt(5) and x+y = 2* sqrt( 2+ _/5) =>x^2 +y^2 = (x+y)^2 - 2xy hope you can simplify further..

OpenStudy (anonymous):

Does it help ?? \[\large x^2 + y^2 = (x+y)^2 - 2xy\]

OpenStudy (anonymous):

nop

OpenStudy (shubhamsrg):

why not ?

OpenStudy (anonymous):

i did nt understand

OpenStudy (anonymous):

\[(x+y)^2 = [(2 + \sqrt{5})^{\frac{1}{2}} +(2 + \sqrt{5})^{\frac{1}{2}} + (2 + \sqrt{5})^{\frac{1}{2}} - (2 + \sqrt{5})^{\frac{1}{2}}]^2\] \[(x+y)^2 = [2(2 + \sqrt{5})^{\frac{1}{2}}]^2 \implies (x+y)^2 = 4(2 + \sqrt{5})\]

OpenStudy (anonymous):

\[((2+\sqrt{5})^{0.5}+(2-\sqrt{5})^{0.5})^{2}=(2+\sqrt{5})^{0.5\times2}+2(2+\sqrt{5})^{0.5}(2-\sqrt{5})^{0.5}+(2-\sqrt{5})^{2}\]

OpenStudy (anonymous):

can you continue after it..?

OpenStudy (anonymous):

\[xy = ((2 + \sqrt{5})^{\frac{1}{2}} + (2 + \sqrt{5})^{\frac{1}{2}})((2 + \sqrt{5})^{\frac{1}{2}} - (2 + \sqrt{5})^{\frac{1}{2}})\] \[xy = (2 + \sqrt{5} - 2 + \sqrt{5}) \implies xy = 2 \sqrt{5}\]

OpenStudy (anonymous):

\[x^2 + y^2 = 8 + 4\sqrt{5} - 4\sqrt{5} = ??\]

OpenStudy (anonymous):

x=\[(2+\sqrt{5})^{1/2} + (2\sqrt{5})^{1/2}\]

OpenStudy (anonymous):

sry.....nt that

OpenStudy (anonymous):

\[x= (2+\sqrt{5})^{1/2} + (2-\sqrt{5})^{1/2} \]

OpenStudy (anonymous):

See: \[(x+y)^2 = [(2 + \sqrt{5})^{\frac{1}{2}} +(2 + \sqrt{5})^{\frac{1}{2}} + (2 + \sqrt{5})^{\frac{1}{2}} - (2 + \sqrt{5})^{\frac{1}{2}}]^2\] \[(x+y)^2 = [2(2 + \sqrt{5})^{\frac{1}{2}}]^2 \implies (x+y)^2 = 4(2 + \sqrt{5})\]

OpenStudy (anonymous):

This is your value for \((x+y)^2\) Now find xy first: \[xy = ((2 + \sqrt{5})^{\frac{1}{2}} + (2 + \sqrt{5})^{\frac{1}{2}})((2 + \sqrt{5})^{\frac{1}{2}} - (2 + \sqrt{5})^{\frac{1}{2}})\] \[xy = (2 + \sqrt{5} - 2 + \sqrt{5}) \implies xy = 2 \sqrt{5}\] So you have now both the values..

OpenStudy (anonymous):

\[\large x^2 + y^2 = 8 + 4\sqrt{5} - 4\sqrt{5} = ??\] Can you solve this now??